If the MPC = 0.75, then the government purchases multiplier is about

Group of answer choices

3

4

1.33

7
Answer #1

I will appreciate if give positive rating

Answer option B) 4

Multiplier= 4

Solution

MPC ( Marginal propensity to consume)=0.75

Multiplier= 1 / 1- MPC

Multiplier= 1 / 1- 0.75

Multiplier= 1/ 0.25

Multiplier= 4